The perimeter of a rectangular brick wall is 68 feet and its area is 280 square feet. What is the length of the longer side of the wall?
Answer by fractalier(2101)

Here
P = 2L + 2W = 68
A = LW = 280
Now divide by two...and get
L + W = 34
W = 34 - L
now substitute
A = LW = L(34 - L) = 280
34L - L^2 = 280
L^2 - 34L + 280 = 0
(L - 14)(L - 20) = 0
L = 14 or L = 20
Thus L = 20 and W = 14